OpenStudy (anonymous):

a triangular wall is made up of one brick on the top layer, 2 bricks on the second then 3 bricks on the third.... if there are 171 bricks altogether, how many layers are there?

OpenStudy (jamesj):

So you want to know: what is an expression for 1 + 2 + 3 + .... + n. Suppose that expression is S(n). Then the question is for what value of n is S(n) = 171 ?

OpenStudy (jamesj):

Do you know the expression for S(n)?

OpenStudy (anonymous):

but I substituted the values into the general equation of Sn for arithmetic sequences and the answer was 300 something which is impossible, could you please show me how to find n?

OpenStudy (anonymous):

So S(n) = n(n+1)/2, hence we want n such that n(n+1)/2 = 171 n^2 + n = 342 Now solve n^2 + n - 342 = 0

OpenStudy (jamesj):

Right. This will give you a negative value and a positive value.

OpenStudy (anonymous):

but the answer is supposed to be 18

OpenStudy (jamesj):

Yes ... that is the positive solution. Obviously the negative solution here is not meaningful.

OpenStudy (anonymous):

oh woops sorry stupid question, haha, I though my answer was either 342 or 341

OpenStudy (jamesj):

ok